OpenStudy (anonymous):

how do you go from x squared +6x+9=7+9 to get (x+3)squared = 16?

OpenStudy (anonymous):

x is a negative number or a decima maybe

OpenStudy (anonymous):

x²+6x+9=7+9=16 x²+6x+9=(x+3)(x+3)=(x+3)²=16 ------------------- this step is what you're referring to?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, when one factors a polynomial, he searches for 2 numbers that, when mutiplied, become c and when added become b (in x²+bx+c). So, we need to find 2 numbers that, when multiplied become 9 and when added become 6.

OpenStudy (anonymous):

...where did you get c and b from?

OpenStudy (anonymous):

Those are variables, just like x and y. I made them up so you could understand to which numbers I'm referring to.

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrquad.htm go to Solve x2 + 6x – 7 = 0 by completing the square. ...i dont get how you go from step 3 to step 4

OpenStudy (anonymous):

What they're doing is, they divide b by 2, and then square that number. They than add that number to both side, so they can complete the square.

OpenStudy (anonymous):

okay i dont get how to "convert the left-hand side to squared form"

OpenStudy (anonymous):

That, I answered in my first answer.

OpenStudy (anonymous):

Oh my god i'm an idiot, thanks.